what are the possible lengths of a rectangle if the width must be 8 inches less the the length and area must be between 100 and 200 square inches? use only the whole numbers for the dimensions
Answer by Boreal(15235)

width=x
length=x+8
area is x(x+8)=x^2+8x
x^2+8x=100
x^2+8x-100=0
x= (1/2) (-8 +/-sqrt (64+400); sqrt 464=21.54. Use positive roots only
x=x=(1/2)13.54=6.77, round to 7 inches
7 x 15=105 in^2.
=======
x^2+8x=200
x^2+8x-200=0
x=(1/2){-8 +/- sqrt (64+800); sqrt 864=29.39
x=(1/2)(21.39=10.70, round down to 10 inches
x+8=18 inches
10 x 18=180 in^2